Four particles of equal masses M move along a circle of radius R under the action of their mutual gravitational attraction. Find the speed of each particle.

Grade:

2 Answers

vikas askiitian expert
13 years ago

net force on any of particle due to remaining three = GM2 (81/2+1)/4R2

centripital force is provided by mutual gravitational force

GM2 (81/2+1)/4R2 = MV2/R

  V = [GM(81/2+1)/4R]

sai raghava rajeev penagamuri
13 years ago

SQUARE ROOT OF GM/R(2(1.414)+1)/4